# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v2

@@ Abstract.
LLM chains enable complex tasks by decomposing work into a sequence of sub-tasks. Crowdsourcing workflows similarly decompose complex tasks into smaller tasks for human crowdworkers. Chains address LLM errors analogously to the way crowdsourcing workflows address human error. To characterize opportunities for LLM chaining, we survey papers across the crowdsourcing and chaining literature to construct a design space for chain development. The design space connects an LLM designer’s objectives to strategies they can use to achieve those objectives, and tactics to implement each strategy. To explore how techniques from crowdsourcing may apply to chaining, we adapt crowdsourcing workflows to implement LLM chains across three case studies: creating a taxonomy, shortening text, and writing a short story. From the design space and our case studies, we identify which techniques transfer from crowdsourcing to LLM chaining and raise implications for future research and development.
Three components: input task, design space, and resulting workflow. Input task is shortening text to specified lengths while preserving meaning. Design space flows through 3 main steps 1) Objectives which are that the outcome quality is verifiable creativity, 2) strategies, which are for response diversity for creativity, and quality control through user revision and verification steps, and 3) tactics of: parallel generation, step to select focus, transform outputs, and evaluate candidates. After design space comes the resulting workflow. This workflow inputs a long text, finds areas to edit, fixes those areas, verifies those fixes, and outputs a shorter text with a length slider. The fix and verify steps have parallel generation.
@@ 1. Introduction
People use Large Language Models (LLMs) for assistance with both single-step tasks, like suggesting synonyms, and more complex multi-step tasks, like writing stories, editing text, and organizing ideas. Despite their widespread adoption, LLMs suffer from quality deficits like hallucinations (Dziri et al., 2022; Nair et al., 2023b; Maynez et al., 2020), brittleness to prompt changes (Gero et al., 2023; Arora et al., 2022; Zamfirescu-Pereira et al., 2023; Nie et al., 2022; Holtzman et al., 2021), and user interventions limited to prompt changes (Wu et al., 2023c; Parameswaran et al., 2023). These deficits are exacerbated for complex tasks, which require multiple steps in which errors can arise and propagate (Dziri et al., 2023). Unfortunately, many common use cases require multiple reasoning steps. For example, the use case of shortening text requires first deciding which portions of the text to edit and then determining how to revise each one (Bernstein et al., 2010).
@@
To tackle complex tasks, recent research has turned to LLM chaining techniques. Chaining decomposes a task into multiple calls to an LLM, in which the output of one call affects the input to the next call (Wu et al., 2021). For example, when shortening text, an LLM chain could identify verbose sentences, edit each one, and propose outputs of variable lengths by composing multiple edits. Chaining has many benefits: this strategy can enable new abilities like long-form story writing (Yang et al., 2022; Mirowski et al., 2023), increase accuracy on logical reasoning questions (Xie et al., 2023; Zhang et al., 2022; Creswell and Shanahan, 2022), and enable greater transparency and debuggability (Huang et al., 2023b; Wu et al., 2022a; Reppert et al., 2023). Although initial evidence suggests that chaining can help mitigate LLM quality deficits, current solutions are imperfect (Nair et al., 2023a; Mirowski et al., 2023; Creswell and Shanahan, 2022; Park et al., 2022; Long et al., 2023). LLM chaining is nascent and challenging, with a seemingly intractable design space.
@@
The challenge of decomposing complex work, however, is not new. In the more established field of crowdsourcing, researchers have faced a similar problem setup. Groups of people (requesters) outsource tasks to other people (crowdworkers) on online platforms. Crowdworkers have their own limitations: approximately of crowdworkers’ outputs are low quality (Bernstein et al., 2010), small changes in instructions affect crowdworker responses (Kittur et al., 2012; McDonnell et al., 2016), and requesters have limited modes of interaction with crowdworkers (Salehi et al., 2017). To compensate for these limitations, crowdsourcing researchers have developed a variety of structured workflows. Crowdsourcing workflows decompose a task into a series of smaller microtasks for crowdworkers to complete. These workflows account for crowdworker limitations by reducing task complexity and leveraging redundancy, aggregation, and explicit verification (Bernstein et al., 2010; Cheng and Bernstein, 2015; Ambati et al., 2012; Zaidan and Callison-Burch, 2011).
@@
Given the parallels between these fields, we investigate how crowdsourcing techniques may be adapted to design LLM chains. Selecting relevant strategies from over a decade of crowdsourcing research is challenging, especially as LLMs and crowdworkers have salient differences (Wu et al., 2023c). For example, crowdworkers may be slow to generate large amounts of text, whereas LLMs may lack common sense and factual grounding, making them more sensitive to instructions and more likely to “hallucinate” information.
@@
To guide LLM chain development, we first construct a design space based on a systematic review of the crowdsourcing workflow and LLM chaining literatures. We analyze papers and perform open coding to identify core design space dimensions. Our design space is organized into three tiers: objectives, strategies, and tactics (Figure 2). The first tier outlines the objectives for which a designer can optimize a workflow. Then, strategies describe approaches to achieve an objective, and tactics are mechanisms that implement strategies.
We then describe how LLM chain designers can navigate this design space. For example, consider a user who wants to shorten input text while still maintaining faithfulness to the original text (Figure 1). This task is analogous to one explored by the Find-Fix-Verify crowdsourcing workflow (Bernstein et al., 2010). Using our design space, the designer would classify their task’s objective as “verifiable creativity”; i.e., a creative task (shortening text) that has to obey some set verifiable constraint (faithfulness). Next, they can browse potential strategies that might help with this objective. They will come across the “response diversity” strategy, which suggests a chain design that redundantly produces multiple shortened outputs that can later be analyzed for faithfulness. To implement this strategy, our design space provides the “parallel generation” tactic by prompting LLMs to produce multiple parallel generations.
@@
To better assess the impact of similarities and differences between crowdsourcing and LLMs, we implement three case studies adapting crowdsourcing workflows to use LLMs. We chose workflows that enable direct manipulation of the output, an instance of the user revision strategy for quality control that has been explored in crowdsourcing but is underexplored in chaining. The case studies implement tasks that vary in their degrees of creativity, adapting Cascade (Chilton et al., 2013) to control the precision of a generated taxonomy, Soylent (Bernstein et al., 2010) to shorten text, and Mechanical Novel (Kim et al., 2017) to control the narrative elements of a fictional story. We evaluate these chains in terms of the quality and controllability of their outputs. For example, compared to zero-shot prompting, chain-generated stories are preferred by crowdsourced raters and the story elements can be more precisely controlled. These case study implementations provide insights into how differences between LLMs and crowdworkers impact chain design to achieve the same objectives. In Mechanical Novel, for instance, we find that the optimal chain requires more granular sub-tasks and extra verification steps than the original crowdsourcing workflow.
Based on our literature survey, design space, and case studies, we identify techniques that do and do not transfer from crowdsourcing and share implications for future work. We find that the objectives, strategies and types of tactics of our design space transfer. We recommend approaches to the user revision strategy and creative work objective applied in crowdsourcing as particularly worth exploration in chaining.
However, to achieve a given strategy, tactic design decisions may not transfer, due to differences in the abilities, strengths, and weaknesses of crowdworkers and LLMs. For example, LLMs can more quickly process and generate fluent text than crowdworkers, different subtask complexities perform best on LLM chains, and LLM chains require more oversight to generate and validate diverse responses. Consequently, future work is needed to study optimal tactic designs, such as tactic selection and implementation, especially within the context of effectively employing certain strategies.
Finally, we explore how the relative speed and reduced cost of LLMs relative to crowdworkers presents an opportunity to create higher quality chains, and we recommend improved tools and methods to augment this design process.
Three tiers of text. The top tiers cover Objectives, strategies, and tactics. Each dimension and subdimension of this top tier is listed, with no other detail.
@@ 2.1. Paper collection
We run two paper collection processes, one for the crowdsourcing literature and another for the LLM chaining literature. We use snowball sampling from the relevant citations in and of these papers, including every citation of the initial AI Chains paper (Wu et al., 2022b). We consider papers available as of August .
@@
Results. Our initial keyword search found papers for workflows and papers for chaining. One author determined whether papers were out of scope given our collection criteria, and any ambiguous cases were discussed among multiple authors. Our final set of papers resulted in papers: for workflows, and for chains.
@@ 2.2. Coding strategy
We conduct a thematic analysis (Braun and Clarke, 2006), following the methodologies of similar design spaces in HCI (Lai et al., 2023; Bae et al., 2022; Shi et al., 2021). First, one author extracted 1) the outcomes and 2) the elements of the workflows from the paper set. We then employed iterative open coding on the extracted data, in which two authors separately coded each data extraction, sorted resulting codes into sub-categories, then discussed these categorizations. In three further rounds of iteration, the same authors refined these categorizations. All authors reviewed the resulting design space. Two authors categorized every paper within the finalized design space.111https://github.com/madeleinegrunde/designingchains
@@ 3. Design Space
Our design space consists of three tiers that cover work in both the crowdsourcing workflow and LLM chaining literatures. First, we cover objectives for which designers may want to optimize: to promote outcome qualities such as accuracy or creativity, while respecting resource constraints. We then delineate strategies by which designers can achieve these objectives: response diversity, quality control mechanisms, and workflow flexibility. Finally, we categorize tactics for implementing these processes, at the levels of workflow architectures, subtasks, and choice of workers. Finally, we demonstrate how to navigate connections between design space tiers, with the objectives guiding design decisions at every level.
We arrived at the tiers through an open coding process. Two of the high level tiers that emerged matched the concept types we extracted from the literature: objectives (outcomes) and tactics (elements). If you were to sketch a workflow on a piece of paper, tactics would be the visible elements in the sketch and the workers who complete each subtask. The process of coding also surfaced a set of abstract concepts that connect the tactics to the objectives. These concepts provide the third intermediate tier: strategies.
@@ 3.1. Objectives in the design space
The objectives of a workflow design consist of achieving a desired outcome quality within resource constraints (Figure 3).
3.1.1. Outcome Quality
Workflows support a range of output quality objectives, from highly accurate to highly creative. To aid discussion, we discretize this spectrum into three categories: tasks that prioritize accuracy, verifiable creativity, or creativity.
@@ 3.1.1. Outcome Quality
Two plots, each of which has several charts, each chart with a Y axis of Prevalence and two bars, one labeled Crowd and one labeled LLM. Plot 1: A set of three bar charts side by side, entitled Outcome Quality. The first chart is labeled Accuracy, with Crowd at 38.2% and LLM at 79.5%. The second chart is labeled Verifiable Creativity, with Crowd at 57.4% and LLM at 28.2%. The third chart is labeled Creativity, with Crowd at 29.4% and LLM at 28.2%. Plot 2: A set of two bar charts side by side, entitled Resource Constraints. The first chart is labeled Expense, with Crowd at 27.9% and LLM at 12.8%. The second chart is labeled Latency, with Crowd at 30.9% and LLM at 2.6%.
Verifiable creativity. In some tasks, there is not one correct answer, but there are restrictions on what outputs are correct. Crowdsourcing papers for these tasks include Cascade (Chilton et al., 2013), Flock (Cheng and Bernstein, 2015), Revolt (Chang et al., 2017), and Storia (Kim and Monroy-Hernandez, 2016). Example tasks include categorizing data (categories must be accurate, but there are multiple options and unclear edge cases) (Bragg et al., 2013; Chilton et al., 2013; André et al., 2013; Chilton et al., 2014; Cheng and Bernstein, 2015; Chang et al., 2017), summarizing text (summaries must stay faithful, but have degrees of freedom in how to write the summary) (Kim and Monroy-Hernandez, 2016; Pavel et al., 2014; Bernstein et al., 2010; Nair et al., 2023a, b), and coding (programs must work, but there are multiple approaches to problem solving) (LaToza et al., 2014; Kulkarni et al., 2014; Wu et al., 2023a; Huang et al., 2023b, a). The goal of workflows for these tasks are to align with expert outputs (Kim et al., 2014; Liu et al., 2018; Chilton et al., 2013; Salehi et al., 2017; Park et al., 2022; Kim et al., 2023b; Fleming et al., 2023) or to enable users to choose among multiple correct possible output options. Using workflows for these tasks can result in an output that is human-understandable (Cheng and Bernstein, 2015), that better represents what would be important for people to know (Salisbury et al., 2017), and that surfaces the uncertainty among the different possible correct options (Chang et al., 2017).

### v3

@@ Abstract.
LLM chains enable complex tasks by decomposing work into a sequence of subtasks. Similarly, the more established techniques of crowdsourcing workflows decompose complex tasks into smaller tasks for human crowdworkers. Chains address LLM errors analogously to the way crowdsourcing workflows address human error. To characterize opportunities for LLM chaining, we survey 107 papers across the crowdsourcing and chaining literature to construct a design space for chain development. The design space covers a designer’s objectives and the tactics used to build workflows. We then surface strategies that mediate how workflows use tactics to achieve objectives. To explore how techniques from crowdsourcing may apply to chaining, we adapt crowdsourcing workflows to implement LLM chains across three case studies: creating a taxonomy, shortening text, and writing a short story. From the design space and our case studies, we identify takeaways for effective chain design and raise implications for future research and development.
Three components: input task, design space, and resulting workflow. Input task is shortening text to specified lengths while preserving meaning. Design space flows through 3 main steps 1) Objectives which are to have a creative response while maintaining accurate actuality, 2) strategies, which are for response diversity, user guidance, and validation, and 3) tactics of: parallel generation, step to select focus, transform outputs, and evaluate candidates. After design space comes the resulting workflow. This workflow inputs a long text, finds areas to edit, fixes those areas, verifies those fixes, and outputs a shorter text with a length slider. The fix and verify steps have parallel generation.
@@ 1. Introduction
People use Large Language Models (LLMs) for assistance with both single-step tasks, like suggesting synonyms, and more complex multi-step tasks, like writing stories, editing text, and organizing ideas. Despite their widespread adoption, LLMs suffer from quality deficits like hallucinations (Dziri et al., 2022; Nair et al., 2023b; Maynez et al., 2020), brittleness to prompt changes (Gero et al., 2023; Arora et al., 2022; Zamfirescu-Pereira et al., 2023; Nie et al., 2022; Holtzman et al., 2021), and user interventions limited to prompting (Wu et al., 2023c; Parameswaran et al., 2023). These deficits are exacerbated for complex tasks, which require multiple steps in which errors can arise and propagate (Dziri et al., 2023). Unfortunately, many common use cases require multiple reasoning steps. For example, the use case of shortening text requires first deciding which portions of the text to edit and then determining how to revise each one (Bernstein et al., 2010).
@@
To tackle complex tasks, recent research has turned to LLM chaining techniques. Chaining decomposes a task into multiple calls to an LLM, in which the output of one call affects the input to the next call (Wu et al., 2021). For example, when shortening text, an LLM chain could identify verbose sentences, edit each one, and propose outputs of variable lengths by composing multiple edits. Chaining has many benefits: this strategy can enable new abilities like long-form story writing (Yang et al., 2022; Mirowski et al., 2023), increase accuracy on logical reasoning questions (Xie et al., 2023; Zhang et al., 2022; Creswell and Shanahan, 2022), and enable greater transparency and debuggability (Huang et al., 2023b; Wu et al., 2022a; Reppert et al., 2023). Although initial evidence suggests that chaining can help mitigate LLM quality deficits, current solutions are imperfect (Nair et al., 2023a; Mirowski et al., 2023; Creswell and Shanahan, 2022; Park et al., 2022; Long et al., 2023; Josifoski et al., 2023). LLM chaining is nascent and challenging, with a seemingly intractable design space.
@@
The challenge of decomposing complex work, however, is not new. In the more established field of crowdsourcing, researchers have faced a similar problem setup. Groups of people (requesters) outsource tasks to other people (crowdworkers) on online platforms. Crowdworkers have their own limitations: approximately 30% of crowdworkers’ outputs are low quality (Bernstein et al., 2010), small changes in instructions affect crowdworker responses (Kittur et al., 2012; McDonnell et al., 2016), and requesters have limited modes of interaction with crowdworkers (Salehi et al., 2017). To compensate for these limitations, crowdsourcing researchers have developed a variety of structured workflows. Crowdsourcing workflows decompose a task into a series of smaller microtasks for crowdworkers to complete. These workflows account for crowdworker limitations by reducing task complexity and leveraging redundancy, aggregation, and explicit validation (Bernstein et al., 2010; Cheng and Bernstein, 2015; Ambati et al., 2012; Zaidan and Callison-Burch, 2011).
@@
Given the parallels between these fields, we investigate how crowdsourcing techniques may be adapted to design LLM chains (Figure 1). Selecting relevant strategies from over a decade of crowdsourcing research is challenging, especially as LLMs and crowdworkers have salient differences (Wu et al., 2023c). For example, crowdworkers may be slow to generate large text amounts, whereas LLMs may lack common sense and factual grounding, making them more sensitive to instructions and likely to “hallucinate” information.
@@
To guide LLM chain development, we first construct a design space based on a systematic review of the crowdsourcing workflow and LLM chaining literatures. We analyze 107 papers and perform open coding to identify core design space dimensions. Our design space is organized into the objectives for which a designer can optimize a workflow (Figure 2) and the tactics used to build the workflow (Figure 3).
From this review, we also surface the strategies used to marshal and combine tactics to achieve objectives. For example, the objective to encourage creativity in outputs can be achieved by strategies that produce diverse responses, allow user guidance, and utilize adaptable architectures. We gather from the literature how tactics have been used in prior work to support strategies. For example, workflows and chains encourage diverse responses with redundant or communicative tactical architectures. Strategies constrain the space and evaluation of tactic choices for both crowds and LLMs (Figure 4).
@@
To better assess the impact of similarities and differences between crowdworkers and LLMs on the implementation of strategies, we implement three case studies adapting crowdsourcing workflows to use LLMs. We chose workflows that enable direct manipulation of the output, an instance of the user guidance strategy for quality control that has been explored in crowdsourcing, but is underexplored in chaining. The case studies implement tasks that vary in their degrees of creativity, adapting Cascade (Chilton et al., 2013) to control the precision of a generated taxonomy, Soylent (Bernstein et al., 2010) to shorten text, and Mechanical Novel (Kim et al., 2017) to control the narrative elements of a fictional story.
We evaluate these chains in terms of the quality and controllability of their outputs. For example, compared to zero-shot prompting, chain-generated stories are preferred by crowdsourced raters and the story elements can be more precisely controlled. These case study implementations provide insights into how differences between LLMs and crowdworkers impact chain design to achieve the same objectives. In Mechanical Novel, for instance, we find that the LLM chain requires more explicit support than the crowdsourcing workflow to create creative and varied outputs.
Informed by both our design space and the three case studies, we discuss how differences between LLMs and crowdworkers impact effective tactic designs for LLM chains. For example, because LLM stochasticity differs from human diversity, tactic choices for LLMs must induce response diversity more explicitly and account for limitations in LLM self-validation. We find that our strategies help focus workflow adaptation and suggest areas of promising future work. Such areas include how the relative speed and reduced cost of LLMs relative to crowdworkers presents an opportunity to create higher quality chains, and we recommend improved tools and methods to augment this design process.
2. Related work
The bulk of our design space contribution is a literature review, so most related work has been discussed inline throughout the paper. Concurrent to this paper, a few other groups investigate the intersection of crowdsourcing and LLMs. One paper explores the impact of incorporating LLMs into crowdsourcing workflows on requester and crowdworker stakeholders (ALLEN et al., 2023). A second considers hybrid Crowd-LLM data annotation (Li, 2024). Another implements three strategies from crowdsourcing using LLMs for data processing tasks (Parameswaran et al., 2023). Closest to our work, the final paper has students implement six crowdsourcing workflows. The paper discusses the effectiveness of these workflows and the impact of the difference between LLM and crowdworker abilities on recommendations for chaining (Wu et al., 2023c). We build on this approach of providing recommendations based on case study implementations. We widen the scope of our recommendations by first integrating a broad survey of LLM chaining and crowdsourcing workflow literature into a design space and then using the case studies to explore open questions in the design space. This broader investigation enables us to build a design space and informs more comprehensive recommendations for navigating design decisions and for orienting future work. These recommendations span from higher-level guidance, such as adhering to the tradeoff between outcome quality and resource constraints, to lower-level suggestions, such as pursuing architecture-based methods for quality validation.
@@ 3.1. Paper collection
We run two paper collection processes, one for the crowdsourcing literature and another for the LLM chaining literature. We use snowball sampling from the relevant citations in and of these papers, including every citation of the initial AI Chains paper (Wu et al., 2022b). We consider papers available as of August 2023.
@@
Results. Our initial keyword search, conducted in August 2023, found 313 papers for workflows and 180 papers for chaining. One author determined whether papers were out of scope given our collection criteria, and any ambiguous cases were discussed among multiple authors. Our final set of papers resulted in 107 papers: 68 for workflows, and 39 for chains.
@@ 3.2. Coding strategy
We conduct a thematic analysis (Braun and Clarke, 2006), following the methodologies of similar design spaces in HCI (Lai et al., 2023; Bae et al., 2022; Shi et al., 2021). First, one author extracted 1) the outcomes (objectives) and 2) the elements (tactics) of the workflows from the paper set. We then employed iterative open coding on the extracted data, in which two authors separately coded each data extraction, sorted resulting codes into sub-categories, then discussed these categorizations. In three further rounds of iteration, the same authors refined these categorizations. All authors reviewed the resulting design space.
@@ 4. Design Space
Our design space has two parts: the objectives span the space of a designer’s goals, and tactics span the space of workflow implementations. These objectives are 1) promoting outcome qualities while respecting resource constraints, 2) increasing the accuracy and/or creativity of outputs, and defining the generality or specificity of the set of tasks that the designer wants the workflow to support. For the workflow tactics, workers complete various subtasks that are linked together into the workflow’s architecture.
Title of ”Objectives, what the designer wants the workflow to accomplish”. Three rows of text each with two objectives, examples, and a description. Row 1. A tradeoff between outcome quality (ex. transparency, no hallucination) and resource constraints (ex. effort, latency, expense). Description: ”Designers must work within resource constraints to improve outcome quality”. Row 2. A spectrum between accuracy (ex. math, question-answering) and creativity (ex. writing, prototyping). Description: Different tasks may require accuracy, creativity, or some combination thereof. Row 3. A spectrum between task specific (ex. only applied to one problem (e.g., translation workflow)) and task general (ex. can be applied to many problems(e.g., map-reduce)). Description: Workflows vary from supporting a specific task to supporting a general set of tasks.
@@ 4.1. Objectives in the design space
The objectives of a workflow design define the designer’s goals within several dimensions. Designers will design for some point along a outcome quality vs resource constraints tradeoff, in a range from accuracy to creativity, and in a range spanning between task specificity and generality. (Figure 2).
4.1.1. Outcome quality vs resource constraints
When working towards outcome qualities, designers must be mindful of resource constraints. Workflows can support quality outputs such as by improving accuracy and/or creativity (Kim et al., 2023b; Creswell and Shanahan, 2022; Lasecki et al., 2013; Reppert et al., 2023; Wu et al., 2021), transparency (Creswell et al., 2022; Creswell and Shanahan, 2022; Wu et al., 2023c; Reppert et al., 2023; Wu et al., 2023a; Kim et al., 2023b), debuggability (Wu et al., 2022a, 2021), and modularity (Huang et al., 2023b; LaToza et al., 2014; Lee et al., 2017; Retelny et al., 2014; Zharikova et al., 2023; Yao et al., 2023; Wu et al., 2023a).
Workflows can also address, although imperfectly, quality concerns. Some quality concerns are induced by the workers, such as hallucinations (Nair et al., 2023a; Mirowski et al., 2023; Creswell and Shanahan, 2022; Nair et al., 2023b; Wu et al., 2023a; Parameswaran et al., 2023) and generative loops (Mirowski et al., 2023) in the case of LLMs, and fatigue (Dai et al., 2013; Alshaibani et al., 2020) in the case of human workers. Others are induced by the workflow itself, such as cascading errors (Wu et al., 2022a, 2021; Xie et al., 2023) and lack of coherency (Bernstein et al., 2010; Yang et al., 2022; Kim et al., 2017; Hahn et al., 2016; Mirowski et al., 2023; Wu et al., 2022a; Hahn et al., 2016). More concerns are induced by the broader context, such as the need to scale (Arora et al., 2022; Kulkarni et al., 2014, 2012), the workflow’s placement within a larger system (Pavel et al., 2014; Christoforaki and Ipeirotis, 2014; Cheng and Bernstein, 2015), unethical use of the workflow (Park et al., 2022), and the impact on human labor (Mirowski et al., 2023; Wu et al., 2023a), sense of ownership (Mirowski et al., 2023) and sense of agency (Kim et al., 2023b; Mirowski et al., 2023). However, mechanisms for improving output quality can be more expensive (Yao et al., 2023; Huffaker et al., 2020), take longer (Chang et al., 2017; Gouravajhala et al., 2018), and require more effort to create (Kittur et al., 2011; LaToza et al., 2014).
Expense. The expense of a workflow refers to paying crowdworkers or paying for an API call to an LLM. Workflows are often more expensive than simpler approaches (e.g., Chain of Thought prompting) (Zharikova et al., 2023; Yao et al., 2023; Wang et al., 2022; Du et al., 2023; Gero et al., 2023; Hahn et al., 2016; Huffaker et al., 2020; Lin et al., 2014a; Dai et al., 2010; Marcus et al., 2011a, b; Nair et al., 2023a). However, the cost is often less than hiring professionals (Ambati et al., 2012; Christoforaki and Ipeirotis, 2014; Zaidan and Callison-Burch, 2011; Liem et al., 2011; Fleming et al., 2023), pretraining models (Kim et al., 2023b; Levine et al., 2022; Yao et al., 2023; Huang et al., 2023b; Cobbe et al., 2021), or completing the task yourself (Zhang et al., 2022; Wang et al., 2023).
Latency. Another constraint is the time for the workflow to complete (Lasecki et al., 2015; Kulkarni et al., 2014; Marcus et al., 2011a, b). Many workflows have lower latency than a single person performing the same task, as work can be parallelized across multiple workers (Teevan et al., 2016; Lasecki et al., 2015; Chilton et al., 2014; Guo et al., 2016; Lee et al., 2017; André et al., 2013; Pietrowicz et al., 2013; Lin et al., 2014a; Gouravajhala et al., 2018; Kulkarni et al., 2014).
Effort. Workflows can help users complete tasks with less effort (Mirowski et al., 2023; Wu et al., 2023a; Kulkarni et al., 2014; Long et al., 2023). However, workflows can require high effort to create (Kulkarni et al., 2012; Hahn et al., 2016; Kittur et al., 2012) and oversee (Wu et al., 2022a; LaToza et al., 2014).
4.1.2. Accuracy vs creativity
One element of output quality of particular importance is if the workflow should support the accuracy or creativity of outputs. Many tasks’ definitions of output quality have elements of both accuracy and creativity.